# EEG2Rep: Enhancing Self-supervised EEG Representation Through Informative Masked Inputs

### v1

@@ 4.2. Competitors and Implementation Setup
We conducted extensive comparisons against five state-of-the-art methods for EEG representation learning. These methods include invariance-based approaches TS-TCC (TS-TCC2021, ),  (TF-C2022, ), , and BIOT (yang2023biot, ) as well as self-prediction methods BENDR (Bendr, ) and MAEEG (maeeg, ). To ensure a fair evaluation, we used publicly available code for the baseline methods. All experiments utilized the PyTorch framework on a system featuring a single Nvidia A5000 GPU (24GB). Model and hyperparameter selection relied on the validation set. Table 2 and Table 3 present five sets of results with varied random seeds, reporting mean and standard deviation values. We adhere to the original VICReg (vicreg2022, ) for setting the hyperparameters $\lambda$, $\mu$, and $\gamma$ in the total loss. For further details on the evaluation metrics, refer to the Appendix B.
4.3. Linear Probing
Table 2 presents the average performance of EEG2Rep along with other state-of-the-art methods over five runs. For each dataset, the number in bold indicates the highest accuracy achieved, while the number underlined represents the second best. The results presented in Table 2 indicate that our model, EEG2Rep, achieves the highest average performance on all EEG tasks.
4.4. Fine Tuning
The comparison encompasses an additional variant of the EEG2Rep model, initialized randomly and denoted as EEG2Rep (Random). This is equivalent to supervised training. The table shows that using pre-trained EEG2Rep leads to an average accuracy improvement of 6% on average compared to EEG2Rep (Random). Significant enhancements are evident in specific datasets, particularly DREAMER, DriverDistraction, and TUEV. In DREAMER and DriverDistraction, EEG2Rep demonstrates AUROC improvement of 5.81% and 2.26%, respectively, when compared to random initialization. Similarly, for TUEV, there is a 6.85% improvement in weighted F1, validating the effectiveness of incorporating informative context input in self-supervised methods for enhanced learning and improved EEG classification. It’s worth noting that in datasets with pronounced subject invariance issues, such as DREAMER, STEW, Driverdistraction, and TUAB, the performance of the supervised models is notably subpar. In some cases, these models even exhibit lower performance compared to models pre-trained self-supervised and utilized for classification through a linear layer. For instance, in DREAMER, linear probed EEG2Rep achieves a 3.84% higher accuracy than its supervised counterpart.
@@ 4.4. Fine Tuning
4.5. Ablation Study
@@
Random Masking 56.97 ($\downarrow$ 10.66)
@@
Masking Ratio/Blocks
@@
Table 4 shows a comparison of linear probing performance when the loss is computed in input-space versus representation space. We hypothesize that a key component of EEG2Rep lies in computing the loss entirely in representation space, enabling the target encoder to generate abstract prediction targets that eliminate irrelevant raw details. As evident from Table 4, predicting in input-space results in a $13.12\%$ degradation in linear probing performance, highlighting the impact of noise and wide range of amplitudes inherent to EEG data.
@@ 5. Conclusion
Appendix A Dataset Overview and Processing Methods
Emotiv
@@ Appendix A Dataset Overview and Processing Methods
Crowdsourced EEG data was collected while participants were engaged in a resting state task, involving periods with eyes open and eyes closed, each lasting 2 minutes. Out of 60 participants, only 12 successfully completed both conditions using 14-channel EPOC+ and EPOC X devices. The data was initially recorded at 2048 Hz and later downsampled to 128 Hz. Raw EEG data for these 12 participants, along with preprocessing, analysis, and visualization scripts, are openly available on the Open Science Framework (OSF) [^5] https://osf.io/9bvgh.
STEW: Simultaneous Task EEG Workload
This dataset comprises raw EEG recordings from 48 participants using a 14-channel Emotiv EPOC headset involved in a multitasking workload experiment utilizing the SIMKAP multitasking test (stew, ). Additionally, the subjects’ baseline brain activity at rest was recorded before the test. The data was captured using the Emotiv Epoc device with a sampling frequency of 128Hz and 14 channels, resulting in 2.5 minutes of EEG recording for each case. Participants were instructed to assess their perceived mental workload after each stage using a rating scale ranging from 1 to 9, and these ratings are available in a separate file. Moreover, this dataset includes binary class labels, considering a workload rating of more than 4 as high and otherwise as low. We utilize these labels for our specific problem. STEW can be accessed upon request through the IEEE DataPort [^6] https://ieee-dataport.org/open-access/stew-simultaneous-task-eeg-workload-dataset.
@@
TUH
@@ Appendix B Details of Experimental Settings
Evaluation Metrics
B.1. Noise Types for Model Robustness

### v2

@@ 4.2. Competitors and Implementation Setup
We conducted extensive comparisons against five state-of-the-art methods for EEG representation learning. These methods include invariance-based approaches TS-TCC (TS-TCC2021, ), TF-C (TF-C2022, ), and BIOT (yang2023biot, ) as well as self-prediction methods BENDR (Bendr, ) and MAEEG (maeeg, ). To ensure a fair evaluation, we used publicly available code for the baseline methods. All experiments utilized the PyTorch framework on a system featuring a single Nvidia A5000 GPU (24GB). Model and hyperparameter selection relied on the validation set. Table 2 and Table 3 present five sets of results with varied random seeds, reporting mean and standard deviation values. We adhere to the original VICReg (vicreg2022, ) for setting the hyperparameters $\lambda$, $\mu$, and $\gamma$ in the total loss. For further details on the evaluation metrics, refer to the Appendix B.
@@
4.3. Linear Probing
Table 2 presents the average performance of EEG2Rep along with other state-of-the-art methods over five runs. For each dataset, the number in bold indicates the highest accuracy achieved, while the number underlined represents the second best (This formatting is consistent across all tables presented in this paper). The results presented in Table 2 indicate that our model, EEG2Rep, achieves the highest average performance on all EEG tasks.
Table 4. EEG2Rep model performance in cross-domain settings on STEW and Crowdsourced datasets.
Models STEW Crowdsourced
ACC AUROC ACC AUROC
Random initialization 70.26$\pm$1.59 70.49$\pm$1.86 91.19$\pm$1.18 91.22$\pm$1.23
In-domain pre-trained 73.60$\pm$1.47 74.40$\pm$1.50 94.13$\pm$2.11 94.13$\pm$2.17
Pre-trained on DREAMER 73.75$\pm$1.95 74.95$\pm$2.77 93.91$\pm$1.78 93.86$\pm$1.80
Pre-trained on DriverDistraction 73.68$\pm$2.17 74.67$\pm$3.06 94.09$\pm$1.95 94.11$\pm$2.02
Pre-trained on All Emotiv 74.11$\pm$2.34 77.38$\pm$2.77 94.05$\pm$1.68 94.07$\pm$1.75
4.4. Fine Tuning
The comparison also includes another version of the EEG2Rep model, initialized randomly and denoted as EEG2Rep (Random). This is equivalent to supervised training. The table shows that using pre-trained EEG2Rep leads to an average accuracy improvement of 6% on average compared to EEG2Rep (Random). Significant enhancements are evident in specific datasets, particularly DREAMER, DriverDistraction, and TUEV. In DREAMER and DriverDistraction, EEG2Rep demonstrates AUROC improvement of 5.81% and 2.26%, respectively, when compared to random initialization. Similarly, for TUEV, there is a 6.85% improvement in weighted F1, validating the effectiveness of incorporating informative context input in self-supervised methods for enhanced learning and improved EEG classification. It’s worth noting that in datasets with pronounced subject invariance issues, such as DREAMER, STEW, Driverdistraction, and TUAB, the performance of the supervised models is notably subpar. In some cases, these models even exhibit lower performance compared to models pre-trained self-supervised and utilized for classification through a linear layer. For instance, in DREAMER, linear probed EEG2Rep achieves a 3.84% higher accuracy than its supervised counterpart.
4.5. Cross-Domain
We evaluated the performance of our EEG2Rep model in a cross-domain setting, where the model is trained on one dataset and tested on another. This approach assesses the model’s ability to generalize across different types of EEG domains and tasks. We utilized two Emotiv datasets for the target tasks: STEW and Crowdsourced, both characterized by limited training samples, making cross-domain evaluation particularly relevant.
As depicted in Table 4, pre-training the model with datasets like DREAMER led to performance improvements for the STEW dataset, where the task is mental workload classification. This improvement can be attributed to the contextual alignment between activities recorded in the DREAMER and STEW datasets. Interestingly, even though the DriverDistraction dataset is not directly related to mental workload classification, the experimental environment and certain distraction classes are similar to those in the STEW dataset. This similarity allowed the EEG2Rep model to benefit from pre-training on the DriverDistraction dataset, as reflected in the performance.
For the Crowdsourced dataset, where the task is eye open/closed classification, the model achieved robust performance when pre-trained on all available Emotiv datasets. However, this performance did not surpass that achieved through in-domain pre-training. This outcome suggests that the alignment of brain activities across datasets is critical for optimal performance and indicates the potential need to utilize a larger number of pre-training datasets to cover various subjects and tasks comprehensively.
@@ 4.5. Cross-Domain
Random Masking 60.19 ($\downarrow$ 7.45)
@@ 4.6. Ablation Study
Masking Ratio
Table 6. Performance comparison between random masking and SSP across different datasets and masking ratios.
Mask Ratio DREAMER Crowdsourced STEW DriverDistraction TUAB TUEV
Random SSP Random SSP Random SSP Random SSP Random SSP Random SSP
10% 52.11 53.02 72.13 76.69 63.13 68.14 64.02 70.41 64.50 72.36 33.01 40.37
25% 53.23 54.73 75.05 77.28 65.12 70.47 65.40 72.94 64.75 73.73 33.14 42.53
50% 52.63 59.89 76.69 81.24 64.97 69.52 64.12 76.12 68.12 76.55 34.59 44.23
75% 51.77 52.03 72.10 76.24 63.27 64.69 63.12 65.03 66.34 70.81 35.28 42.19
90% 51.51 51.95 70.77 73.44 62.05 63.70 62.06 64.91 66.34 70.12 34.14 38.72
Average 52.25 54.32 73.35 76.98 63.71 67.30 63.74 69.88 66.01 72.71 34.03 41.61
The masking ratio during self-supervised pretraining is crucial as it determines both the difficulty of the self-prediction task and the quality of the learned representations. Therefore, identifying an optimal masking ratio is essential for effective representation learning from EEG data. Following the recent study on the effects of masking on representation learning (kong2023understanding, ), we pre-train our model using five different mask rates [10%, 25%, 50%, 75%, 90%] with our SSP method and random masking to evaluate the quality of representation learning. As shown in Table 6, several interesting findings emerged from the results. We observed that conservative masking (10%) leads to low performance, regardless of the masking strategy. However, as the masking ratio increases, the performance of the model improves, indicating that higher masking can result in more high-level representation learning. Nevertheless, performance degradation occurs with overly aggressive masking (90%), suggesting that representation does not become monotonically more high-level with increasing masking aggressiveness. In other words, overly aggressive masking also leads to low-level representations, similar to conservative masking. Moderate and high masking ratios yield the best results for both random and SSP masking strategies.
As shown in Table 6, our SSP masking strategy consistently achieves higher performance than random masking across various masking ratios. A key distinction between our proposed masking strategy in EEG2Rep and random masking lies in the arrangement of visible patches. In random masking, there is a possibility that visible patches are not consecutive, whereas in EEG2Rep, visible patches (referred to as preserved patches) are ensured to remain contiguous. Additionally, during the reconstruction of masked patches near the boundary between masked and unmasked regions, the model uses nearby visible patches to interpolate, thereby capturing low-level information, which resonates with the empirical observations in (mae2022, ; kong2023understanding, ). Our proposed masking strategy is less susceptible to such phenomena compared to random masking, as it ensures that masked tokens remain contiguous. The high number of boundaries in random masking may result in less contextualized learning. Our empirical results further confirm that our proposed method yields higher performance compared to random masking, even with aggressive and conservative masking.
Masking Blocks
@@
Table 5 shows a comparison of linear probing performance when the loss is computed in input-space versus representation space. We hypothesize that a key component of EEG2Rep lies in computing the loss entirely in the representation space, enabling the target encoder to generate abstract prediction targets that eliminate irrelevant raw details. As evident from Table 5, predicting in input-space leads to a $13.12\%$ degradation in linear probing performance, highlighting the impact of noise and wide range of amplitudes inherent to EEG data.
@@ 5. Conclusion
Appendix A Dataset Overview and Processing
A.1. Emotiv
@@
Crowdsourced EEG data was collected while participants were engaged in a resting state task, involving periods with eyes open and eyes closed, each lasting 2 minutes. Out of 60 participants, only 13 successfully completed both conditions using 14-channel EPOC+, EPOC X, and EPOC devices. The data was initially recorded at 2048 Hz and later downsampled to 128 Hz. Raw EEG data for these 13 participants, along with preprocessing, analysis, and visualization scripts, are openly available on the Open Science Framework (OSF) [^5] https://osf.io/9bvgh.
Simultaneous Task EEG Workload (STEW)
STEW dataset comprises raw EEG recordings from 48 participants using a 14-channel Emotiv EPOC headset involved in a multitasking workload experiment utilizing the SIMKAP multitasking test (stew, ). Additionally, the subjects’ baseline brain activity at rest was recorded before the test. The data was captured using the Emotiv Epoc device with a sampling frequency of 128Hz and 14 channels, resulting in 2.5 minutes of EEG recording for each case. Participants were instructed to assess their perceived mental workload after each stage using a rating scale ranging from 1 to 9, and these ratings are available in a separate file. Moreover, this dataset includes binary class labels, considering a workload rating of more than 4 as high and otherwise as low. We utilize these labels for our specific problem. STEW can be accessed upon request through the IEEE DataPort [^6] https://ieee-dataport.org/open-access/stew-simultaneous-task-eeg-workload-dataset.
@@
A.2. Temple University Hospital (TUH) EEG Corpus
TUH Abnormal EEG Corpus (TUAB)
The TUH Abnormal EEG Corpus (TUAB) is a subset of the Temple University Hospital (TUH) EEG Corpus, which is one of the largest publicly available collections of clinical EEG data. The TUAB specifically focuses on EEG recordings labeled as abnormal, making it a valuable resource for studies on neurological disorders, brain function anomalies, and the development of diagnostic tools (TUAB, ).
TUH EEG Events (TUEV)
The TUH EEG Events Corpus (TUEV) contains annotations of EEG segments classified into six different categories: spike and sharp wave, generalized periodic epileptiform discharges, periodic lateralized epileptiform discharges, eye movement, artifact, and background (TUEV, ).
Acquisition and Preprocessing
@@ Appendix B Details of Experimental Settings
B.1. Parameter Setting
In our experiment, the EEG2Rep model employed one depth-wise and two spatial-wise convolution layers for input embedding, each with 16 filters. During training, a batch size of 256 was used, and we utilized the Adam optimization algorithm (kingma2014adam, ). To prevent overfitting, we implemented an early stopping method based on the validation loss. The model was pre-trained for 500 epochs, after which logistic regression was applied to the representations for linear probing. Similar to the transformer-based model for multivariate time series classification (TST) (TST2021, ) and the default transformer block (attention2017, ), in our experiments, we employed eight attention heads in both the context and target encoder to capture diverse features from the EEG. The transformer encoding dimension was set to $d_{e}=16$, and the feed-forward network (FFN) in the transformer block expanded the input size by a factor of 4 before projecting it back to its original size. For the learning rate, we start with $1\times 10^{-3}$ and use a cosine learning rate scheduler to adjust it over time (data2vec2022, ).
B.2. Noise Types for Model Robustness